Assume that the average annual salary for a worker in the United States is $36,000 and that the annual salaries for Americans are normally distributed with a standard deviation equal to $7,000. Find the following:

(A) What percentage of Americans earn below $26,000?
(B) What percentage of Americans earn above $39,000?

Answer by stanbon(75887)

Assume that the average annual salary for a worker in the United States is $36,000 and that the annual salaries for Americans are normally distributed with a standard deviation equal to $7,000. Find the following:
(A) What percentage of Americans earn below $26,000?
z = (26000-36000)/7000 = -1,4286
P(x < 7000) = P(z < -1.4286) = 0.0766
===================================
(B) What percentage of Americans earn above $39,000?
z = (39000-36000)/7000 = 0.4286
P(x > 39000) = P(z > 0.4286) = 0.3341
=========================
Cheers,
Stan H.